# Lifelong and Universal Machine Learning Potentials for Chemical Reaction Network Explorations

### v1

@@ 4.6 Lifelong Machine Learning Potentials for Exploration of Chemical Reaction Networks
Figure 8: RMSEs of (a) energies $\overline{E}$ and (b) atomic force components $\overline{F}_{\alpha,n}$ evaluated for all data sets employing lMLP ensembles trained on different numbers of training data sets $n_{\mathrm{trained}}$ (applying lADS scheme (3)). Each training data set is a random subset of the respective data set, i.e., training data are contained in the RMSE evaluation. Lines are solid up to the last trained data set of the respective lMLP. The extrapolation to the data set to be trained next is shown as dashed line, while further extrapolations are connected by dotted lines.
@@
Finally, the accuracy of an lMLP ensemble is analyzed in each training stage for previously trained and incoming data. Figures 8 (a) and (b) show the errors for each number of trained data sets $n_{\mathrm{trained}}$ in our simulation of a rolling CRN exploration with data sets $n_{\mathrm{data}}=\{1,...,16\}$. In general, the lMLP shows in every training stage very good accuracy for the data on which it was already trained. The energy accuracy is relatively constant for all trained data sets. The force RMSE increases slightly with increasing $n_{\mathrm{data}}$. However, some variation is expected because the complexity of the training structures increases with $n_{\mathrm{data}}$ while the model architecture stays constant. Hence, the model could initially be too complex for the data and finally be affected by capacity issues.
@@
We point out that the lMLP accuracy for the initial data remains almost constant with an increasing number of training stages. Only for the forces, a slight increase in error is notable. However, as the error also slightly increases for additionally trained data, the reason can be a capacity issue due to the constant model architecture. Consequently, the small subset of training data chosen by lADS is sufficient to retain previous expertise. Extrapolations to data to be trained result in higher errors, highlighting the efficiency of training. The more data sets are in between the last trained set and the evaluated set, the higher the error, since the ordering of the CRN data leads to an increasing difference in the data according to the progress of the chemical process. We note that the errors of the extrapolated predictions are still lower than the errors of the uMLP predictions (see Figure 3). Hence, the lMLP can efficiently and continuously learn additional data, while previous expertise is kept. Therefore, the lMLP approach is applicable for a rolling exploration. To produce efficient and accurate models, even when training is continued for a large number of additional data points, we expect algorithms to come into play that adjust (grow and shrink) the model architecture during training [116].
[…]
@@
Table 2: Mean RMSEs of individual HDNNP ensemble members for final training data $E^{\mathrm{train(final)}}$ and $F_{\alpha,n}^{\mathrm{train(final)}}$, i.e., those data which are not disregarded in the last epoch, and test data $E^{\mathrm{test}}$ and $F_{\alpha,n}^{\mathrm{test}}$. The given errors are the respective standard deviations. The RMSEs of the HDNNP ensemble are evaluated on all data $\overline{E}$ and $\overline{F}_{\alpha,n}$.
@@
The results in Figures 9 (a) and (b) include the evaluation of training data (in contrast to Figures 3 to 6). For a fairer comparison with the uMLP results, pure test data need to be employed which are available for the individual HDNNPs of the lMLP ensemble. Table 2 shows that the mean RMSE of the test energies also satisfies the chemical accuracy criterion. The mean RMSE of the test atomic force components is close to the target threshold. Since the ensemble results are, in general, better than the individual ensemble member predictions, the force RMSE is sufficiently small. We note that the RMSE of the final training data can be larger than the test RMSE because about two thirds of well represented training data were sorted out in the final epoch, while the test data were randomly chosen before training and remained unchanged.
@@ 5 Conclusions
In this work, we evaluated the applicability of universal machine learning potentials (uMLPs) and lifelong machine learning potentials (lMLPs) to drive the exploration of chemical reaction networks (CRNs). While the generalization capability of current uMLP parametrizations is not yet high enough for search trials in CRN explorations, lMLPs offer a promising alternative. An lMLP is a representation of the potential energy surface for arbitrary systems with uncertainty quantification that can be fine-tuned and extended in a rolling fashion. Hence, it unites accuracy, efficiency, and flexibility. In combination with a $\Delta$-learning approach, our lMLP could reach chemical accuracy for the given $\text{HCN}+\text{H}_{2}\text{O}$ CRN data stream.
We proposed a modified lifelong adaptive data selection (lADS) algorithm to improve the continual learning performance of an lMLP. With this algorithm, an lMLP can handle conformation space extensions efficiently. The resulting accuracy is similar to that obtained by learning all data from the start, which is a much easier learning case but not feasible in applications with rolling data influx. The training data required to counteract forgetting can be reduced by lADS to a third in a reliable and stable way, while the accuracy of previous test data remains high. The latter is proven to be true even after adding data sets for ${}^{\text{th}}$ time, whereby each data set contained the same number of structures as the initial data set. Consequently, this evaluation of continual learning performance is significantly beyond our initial proof for one single addition of data [47]. Moreover, our results confirm that continual learning is able to take advantage of already acquired expertise to improve the final accuracy compared to training from scratch for the same number of epochs. Furthermore, we found that lADS and the continual learning features of the CoRe optimizer can improve the final accuracy not only in continual learning but also in learning stationary data.
Consequently, this work can be considered a proof of principle that lMLPs have all attributes to explore CRNs in a rolling fashion. Hence, lMLPs can be reliably applied on-the-fly during an exploration, where the CRN will be generated directly with an lMLP instead of DFT as in this work. Based on the uncertainty quantification (accessible in an MLP ensemble approach) it is then possible to decide on where additional DFT data needs to be generated for the refinement of the lMLP in subsequent lifelong learning epochs. To start such a process from a reasonable initial lMLP, the lMLP can be pre-trained on a large and diverse data set such as one of those employed for uMLPs. Naturally, the demand for such initial training events becomes rarer with increasing number of lMLP applications and, potentially, by community efforts to build generally applicable lMLPs (i.e., lMLPs that inherit universal features from diverse applications). Subsequently, continual learning can be applied to train on unknown structures occurring in the actual simulations of interest. In this way, the approach is similar to transfer learning or fine-tuning of foundation models on system-specific data, but continual learning harbors the advantage that learning can be continued for much more than one iteration.
@@
The lMLP software will be made available on GitHub (https://github.com/ReiherGroup/lMLP) and PyPI (https://pypi.org/project/lmlp).
The SCINE Parrot software module will be made available on GitHub (https://github.com/qcscine/parrot) and PyPI (https://pypi.org/project/scine-parrot).

### v2

@@ 4.5 Lifelong Machine Learning Potentials for Exploration of Chemical Reaction Networks
Figure 8: RMSEs of (a) energies $\overline{E}$ and (b) atomic force components $\overline{F}_{\alpha,n}$ evaluated with lMLP ensembles at different training stages. The total CRN benchmark data was split into $16$ chronologically ordered sets of equal size, with indices $n_{\mathrm{data}}=\{1,...,16\}$. The lMLP ensembles were trained on the sets $n_{\mathrm{data}}=1$ to $n_{\mathrm{trained}}$. The training data set of each lMLP ensemble member was a random subset of the respective sets. The RMSE evaluation of the lMLP ensembles was based on all data of the respective set $n_{\mathrm{data}}$ and hence contained trained data. Lines are solid up to the last trained data set of the respective lMLP ensemble. The extrapolation to the data set to be trained next is shown as dashed line, while further extrapolations are connected by dotted lines. In this figure and in Figure 9, lines are shown to guide the eye, but only values at integer numbers of $n_{\mathrm{data}}$ are meaningful.
@@
Finally, the accuracy of an lMLP ensemble is analyzed in each training stage for previously trained and incoming data. Figures 8 (a) and (b) show the errors for each number of trained data sets $n_{\mathrm{trained}}$ in our simulation of a rolling CRN exploration with data sets $n_{\mathrm{data}}=\{1,...,16\}$. In general, the lMLP shows in every training stage very good accuracy for the data on which it was already trained. The energy accuracy is relatively constant for all trained data sets. The force RMSE increases slightly with increasing $n_{\mathrm{data}}$. However, some variation is expected because, on the one hand, more data needs to be well represented and the complexity of the structures increases with $n_{\mathrm{data}}$, while, on the other hand, the model architecture stays constant. Hence, the model could initially be too complex for the data and finally be affected by capacity issues.
@@
We point out that the lMLP accuracy for the initial data remains almost constant with an increasing number of training stages. Only for the forces, a slight increase in error is notable. However, as the error also slightly increases for additionally trained data, the reason can be a capacity issue due to the constant model architecture. Consequently, the small subset of training data chosen by lADS is sufficient to retain previous expertise. Predictions of data to be trained in the upcoming training stages result in higher errors, highlighting the efficacy of training. The more data sets are in between the last trained set and the evaluated untrained set, the higher the error. The reason is the chronological ordering of the CRN data that leads to an increasing difference in the data, since the chemical reaction process progresses. Hence, the lMLP can efficiently and continuously learn additional data, while previous expertise is kept. Therefore, the lMLP approach is applicable for a rolling exploration. To produce efficient and accurate models, even when training is continued for a large number of additional data points, we expect algorithms to come into play that adjust (grow and shrink) the model architecture during training [101].
To point out the benefit of the continual learning algorithms, we can compare Figures 8 (a) and (b) to Figures S2 (a) and (b) in the Supporting Information. The latter figures show the results for the same training task, but without applying rehearsal of training data, lifelong adaptive data selection, and the stability plasticity balance of the CoRe optimizer. Hence, we only trained on the added data starting from the previously obtained model weight parameters. In this case, the MLP ensembles also show very good accuracy for the respective last trained data set. However, the error increases a lot, the more data sets are in between the last trained set and the evaluated set. This trend does not only apply to upcoming training data sets, but also to data sets that were already trained. For example, the MLP ensemble yields the highest RMSE value for the first trained data set after training on all data sets compared to any previous training stage. The value of this RMSE is even similar to the RMSE value on the last data set of the MLP ensemble that was only trained on the first data set. Therefore, catastrophic forgetting occurs in this training task when we do not apply continual learning. If all previous training data are used in every training, we will also avoid catastrophic forgetting. However, this approach will lead to much higher computational demand compared to continual learning. We note that a sequence of several fine-tunings of a foundation model in a row can lead to similar forgetting if continual learning is not applied. Hence, even the expertise on previously fine-tuned data may be forgotten after a sequence of model fine-tunings.
Figure 9: RMSEs of (a) energies $\overline{E}$ and (b) atomic force components $\overline{F}_{\alpha,n}$ evaluated for all data sets $n_{\mathrm{data}}=\{1,...,16\}$ employing an lMLP ensemble trained on PBE data and an lMLP ensemble trained on $\text{PBE}-\text{GFN2}$ data. The RMSE interval shaded in green represents the accuracy aimed for.
To enable reliable kinetic modeling based on the CRN energies, we require at least chemical accuracy, i.e., $1\,\mathrm{kcal\,mol}^{-1}=4.184\,\mathrm{kJ\,mol}^{-1}$, or better. Since the benchmark data contains systems with up to 12 atoms, a minimum energy accuracy of about $3.614\,\mathrm{meV\,atom}^{-1}$ must be the target. Figure 9 (a) reveals that the lMLP ensemble accuracy is below this target threshold. Similarly, the MLP accuracy target of $100\,\mathrm{meV}\,\text{\AA }^{-1}$ for the atomic force components [69] is satisfied (Figure 9 (b)). In addition, Figures 9 (a) and (b) show the advantage of the $\Delta$-learning approach. If an lMLP is trained in the same way on pure PBE data, the resulting energy and force RMSEs will be approximately twice as large as those of the lMLP trained on $\text{PBE}-\text{GFN2}$ data. This reduction in error is required to yield chemical accuracy for this CRN data set with the given MLP base method.
@@
Table 2: RMSEs of individual HDNNP ensemble members and the HDNNP ensemble (trained on $\text{PBE}-\text{GFN2}$ data). For the individual HDNNP ensemble members, the mean and standard deviation of every RMSE are given for the final training data $E^{\mathrm{train(final)}}$ and $F_{\alpha,n}^{\mathrm{train(final)}}$, i.e., those data which were not disregarded in the last epoch, and test data $E^{\mathrm{test}}$ and $F_{\alpha,n}^{\mathrm{test}}$. For the HDNNP ensemble, the RMSEs were evaluated on all data $\overline{E}$ and $\overline{F}_{\alpha,n}$.
@@
The results in Figures 8 and 9 include the evaluation of training data (in contrast to Figures 4 to 6). For an unbiased evaluation, pure test data need to be employed which are available for the individual HDNNPs of the lMLP ensemble. Table 2 shows that the mean RMSE of the test energies also satisfies the chemical accuracy criterion. The mean RMSE of the test atomic force components is close to the target accuracy. Since the ensemble results are, in general, better than the individual ensemble member predictions, the force RMSE is sufficiently small. We note that the RMSE of the final training data set can be larger than the test RMSE because about two thirds of well represented data were sorted out from the training data set during training, while the test data set was randomly chosen before training and remained unchanged.
@@ 5 Conclusions
In this work, we evaluated the applicability of lifelong machine learning potentials (lMLPs) to drive the exploration of chemical reaction networks (CRNs). An lMLP is a representation of the potential energy surface for arbitrary systems with uncertainty quantification that can be fine-tuned and extended in a rolling fashion. Hence, it unites accuracy, efficiency, and flexibility. In combination with a $\Delta$-learning approach, our lMLP can reach chemical accuracy for the given $\text{HCN}+\text{H}_{2}\text{O}$ CRN data stream.
[…]
Consequently, this work can be considered a proof of principle that lMLPs have all attributes to explore CRNs in a rolling fashion. Hence, lMLPs can be reliably applied on-the-fly during an exploration, where the CRN will be generated directly with an lMLP (instead of DFT as in this work). Based on the uncertainty quantification (accessible in an MLP ensemble approach) it is then possible to decide on where additional DFT data need to be generated for the refinement of the lMLP in subsequent lifelong learning epochs. Uncertain results can then be either directly replaced by the DFT data or recalculated by the improved lMLP afterwards.
To start such a process from a reasonable initial lMLP, the lMLP can be pre-trained on a large and diverse data set such as one of those employed for foundation models or universal MLPs. We note that the lMLP concept can be applied to other MLP methods than HDNNPs as well, enabling the usage of an already pre-trained foundation model as initial lMLP parametrization. With the increasing size of the data sets used for pre-training of foundation models [102] as well as the increasing number of diverse lMLP applications, the demand for such initial training events becomes rarer. By community efforts, we can head towards lMLPs that are generally applicable out-of-the-box. However, due to the enormous size of the chemical space, this is a long way to go. Therefore, continual learning is required to train efficiently on unknown or insufficiently represented structures occurring in the actual simulations of interest. In this way, the approach is similar to transfer learning or fine-tuning of foundation models on system-specific data, but continual learning harbors the advantage that learning can be continued for much more than one iteration.
@@
The lMLP software will be available on GitHub (https://github.com/ReiherGroup/lMLP) and PyPI (https://pypi.org/project/lmlp.
The SCINE Parrot software module will be available on GitHub (https://github.com/qcscine/parrot) and PyPI (https://pypi.org/project/scine-parrot).
@@ Acknowledgment
Supporting Information
Summary of the HDNNP method, description of eeACSF descriptors, tables of eeACSF parameter values, tables of reference element energy values, table of reference data composition (in terms of data origin), evolution of test RMSEs during lMLP training, and performance monitoring of incremental CRN data training without lifelong learning (PDF file).